# Chain-of-Table: Evolving Tables in the Reasoning Chain for Table Understanding

### 09 Jan 2024 v1

@@ 1 Introduction
On the other hand, inference on tables typically involves a series of intermediate reasoning steps and each of them aligns with specific tabular operations. We propose Chain-of-Table, where we conduct step-by-step reasoning as step-by-step tabular operations to form a chain of tables. The tables in the chain are the transformed tables by the tabular operations, representing the intermediate reasoning results. This procedure resembles the thought of reasoning in Chain-of-Thought (Wei et al., 2022). Specifically, we define a set of table operations, such as adding columns, selecting rows, grouping, and more, which are commonly-used in SQL and DataFrame development  (Pönighaus, 1995; Shi et al., 2020; Katsogiannis-Meimarakis and Koutrika, 2023). We then prompt LLMs to conduct step-by-step reasoning. In each step, the LLM dynamically generates an operation as the next step along with its required arguments, and then we execute the operation on the table programmatically. This operation can either enrich the table by adding detailed intermediate results or condense it by removing irrelevant information. Intuitively, visualizing the intermediate results is essential for reaching correct predictions. We feed the transformed table back for the next step. This iterative process continues until an ending state is achieved. We argue that the tables obtained during the reasoning steps are better structured representations of the intermediate thoughts than free-form text. Finally, the Chain-of-Table reasoning results in tables from which it is easier for LLMs to derive a final answer to the question.
We validate Chain-of-Table with three tabular benchmarks to evaluate table-based reasoning: WikiTQ (Pasupat and Liang, 2015), TabFact (Chen et al., 2019), and FeTaQA (Nan et al., 2022). We conduct our experiments using proprietary PaLM 2 (Anil et al., 2023) and GPT-3.5 (Brown et al., 2020; OpenAI, 2023), and open-sourced LLaMA 2 (Touvron et al., 2023), to demonstrate that our proposed method Chain-of-Table is able to generalize to various LLM options. We summarize our contribution as follows:
@@ 2 Related Work
LLMs can learn from a few samples as prompts through in-context learning. This strategy is widely used to give models additional instructions to better solve downstream tasks. Chain-of-Thought (CoT) (Wei et al., 2022) proposes to generate reasoning steps before answering instead of directly generating an end-to-end answer. Following CoT, Least-to-Most (Zhou et al., 2022) and DecomP (Khot et al., 2022) propose to break down the question into subproblems in the reasoning chain. During reasoning, the latter steps are aware of the previous ones. Such iterative chains with task decomposition further improve the results on complex problems by leveraging the intermediate results from solving subproblems. Jin and Lu (2023) enhances CoT through a table-filling procedure, with a primary focus on text-based tasks where the input and output are in textual format. However, the line of works following CoT is not specifically designed for tabular data. As reported in Chen (2023), large language models with these generic reasoning methods can achieve decent results, but there are still gaps between these methods and those specialized for table scenarios (Cheng et al., 2022; Ye et al., 2023). We propose Chain-of-Table to fill the gap by directly incorporating intermediate tables from tabular operations as a proxy of intermediate thoughts.
To better solve table-based tasks with LLMs, researchers go beyond general text and resort to using external tools. Chen et al. (2022); Gao et al. (2023) propose solving reasoning tasks by generating Python programs, which are then executed using the Python interpreter. This approach greatly improves the performance of arithmetic reasoning. In the scenario of table understanding, Text-to-SQL with LLMs (Rajkumar et al., 2022) is a straightforward application of this idea. To further push the limits of programs, Binder (Cheng et al., 2022) generates SQL or Python programs and extends their capabilities by calling LLMs as APIs in the programs. LEVER (Ni et al., 2023) also proposes solving the table-based tasks with programs but with the additional step of verifying the generated programs with their execution results. However, in all of these program-aided methods, the assistant programs still fall short in solving difficult cases that involve complex tables. These limitations are primarily due to the constraints of the single-pass generation process, where the LLMs lack the capability to modify the table in response to a specific question, requiring them to perform reasoning over a static table. Our method, on the contrary, is a multi-step reasoning framework that conducts tabular reasoning step by step. It transforms the tables to be tailored to the given question.
@@ 1 Function Chain-of-Table ($T$, $Q$):
      /* Initialize the operation chain chain with [B] and $\phi$, where */
      /* [B] is the beginning tag and $\phi$ means it requires no arguments */
@@
             $\triangleright$ operation chain
@@ 4 Experiments
We evaluate the proposed Chain-of-Table on three public table understanding benchmarks: WikiTQ (Pasupat and Liang, 2015), FeTaQA (Nan et al., 2022), and TabFact (Chen et al., 2019). WikiTQ and FeTaQA are datasets focused on table-based question answering. They require complex tabular reasoning over the provided table to answer questions. WikiTQ typically requires short text span answers, whereas FeTaQA demands longer, free-form responses. TabFact, on the other hand, is a table-based binary fact verification benchmark. The task is to ascertain the truthfulness of a given statement based on the table. For WikiTQ evaluation, we use the official denotation accuracy (Pasupat and Liang, 2015), and for TabFact, we employ the binary classification accuracy. Given the nature of FeTaQA, which involves comparing predictions with longer target texts, we utilize BLEU (Papineni et al., 2002), ROUGE-1, ROUGE-2, and ROUGE-L (Lin, 2004) for assessment.
In our experiments, we use PaLM 2-S [^1] https://cloud.google.com/vertex-ai/docs/generative-ai/learn/generative-ai-studio, GPT 3.5 (turbo-16k-0613) [^2] http://openai.com/api/, and LLaMA 2 (Llama-2-17B-chat) [^3] https://ai.meta.com/llama/ as the backbone LLMs. We incorporate few-shot demo samples from the training set into the prompts to perform in-context learning. Examples of these prompts can be found in Appendix E. Details regarding the LLM inference parameters and the number of demonstration samples used are provided in Appendix C.
@@ 4.1 Baselines
End-to-End QA guides the LLM to directly produce the answer when provided with a table and a question as input prompts. Few-Shot QA operates similarly, but it includes few-shot examples of (Table, Question, Answer) triplets in the prompt, as detailed in Brown et al. (2020). We select these examples from the training set, and the model also outputs the answer directly. Chain-of-Thought (Wei et al., 2022) prompts the LLM to articulate its reasoning process in text format before delivering the question. Examples of prompts for each of these baselines are available in Appendix F.
@@ 4.2 Results
We compare Chain-of-Table with generic reasoning methods and program-aided reasoning methods on three datasets: WikiTQ, TabFact, and FeTaQA. The results on WikiTQ and TabFact are presented in Table 1444Dater Ye et al. (2023) with OpenAI Codex LLM achieves 65.9% and 85.6% accuracy on WikiTQ and TabFact, respectively. It also achieves 27.96 in BLEU, 0.62 in ROUGE-1, 0.40 in ROUGE-2, and 0.52 in ROUGE-L on FeTaQA. However, because Codex is no longer publicly available, we do not compare Chain-of-Table with Dater with Codex.. We have additional results on FeTaQA in Appendix B. We follow the previous works and report the performance using the official evaluation pipeline.
@@
We also observe that our proposed Chain-of-Table is effective across all backbone models experimented, while other competing methods, such as Binder, perform better on larger LLMs but its performance decreases with smaller LLaMA 2 (Llama-2-17B-chat).
@@
4.3 Performance Analysis under Different Operation Chain Lengths
Dataset Length of operation chain
1 2 3 4 5
WikiTQ 95 1308 1481 1084 341
TabFact 4 547 732 517 223
Figure 3: Performance of Chain-of-Thought, Dater, and the proposed Chain-of-Table on WikiTQ for questions that require an operation chain of varying lengths. Our proposed atomic operations allow our method to dynamically transform the input table through multiple reasoning iterations. This significantly improves performance over generic and program-aided reasoning counterparts.
@@ 4.3 Performance Analysis under Different Operation Chain Lengths
4.4 Performance Analysis under Different Table Sizes
Large tables present significant challenges to LLMs since LLMs often struggle to interpret and integrate contexts in long input prompts (Liu et al., 2023a; Ye et al., 2023). To assess the performance on tables of various sizes, we categorize the input tables from WikiTQ into 3 groups based on token count: small ($<$2000 tokens), medium (2000 to 4000 tokens) and large ($>$4000 tokens). We then compare Chain-of-Table with Dater (Ye et al., 2023) and Binder (Cheng et al., 2022), the two latest and strongest baselines, as representative methods. Detailed performance results are presented in Table 2.
4.5 Efficiency Analysis of Chain-of-Table
We analyze the efficiency of Chain-of-Table by evaluating the number of required generated samples. We compare Chain-of-Table with Binder (Cheng et al., 2022) and Dater (Ye et al., 2023), the two latest and most competitive baseline method. The analysis results on WikiTQ are presented in Table 4. Binder generates Neural-SQL queries, requiring 50 samples for self-consistent results. Dater involves multiple steps, such as decomposing the tables and generating cloze queries for the questions. In each step, Dater also employs self-consistency to improve accuracy, leading to a high number of required generated samples. For a detailed description of these frameworks, please refer to the corresponding papers, Ye et al. (2023) and Cheng et al. (2022).
Prompting Total # of # of generated
generated samples in
samples each steps
Binder (Cheng et al., 2022) 50 Generate Neural-SQL: 50
Dater (Ye et al., 2023) 100 Decompose Table: 40;
Generate Cloze: 20;
Generate SQL: 20;
Query: 20
Chain-of-Table (ours) $\leq$25 DynamicPlan: $\leq$5;
GenerateArgs: $\leq$19;
Query: 1
@@ A.2 Ablation Study
To demonstrate the effectiveness of our proposed atomic operations, we perform an ablation study by creating five leave-one-out variants of our method, each of which removes one of the pre-defined operations from the pre-defined operation pool. For example, w/o f_add_column() means f_add_column() is removed from the operation pool. As a result, the LLM is only able to plan from the remaining four operations (f_select_column, f_select_row, f_group_by, f_sort_by) to construct an operation chain. We report the results of the ablation study in Table 5.
@@ Appendix B Experiments of Chain-of-Table on FeTaQA
Table 6 shows that Chain-of-Table also improves the performance of free-form question answering on FeTaQA across all metrics, whereas Dater (Ye et al., 2023) fails to improve the ROUGE scores compared with End-to-End QA. We also observe the marginal improvement of Chain-of-Table compared with the baseline methods. We attribute this to the nature of the n-gram text similarity metrics of ROUGE-1/2/L (Lin, 2004). As discussed in Maynez et al. (2023); Dhingra et al. (2019), these metrics are known to be insensitive to capturing improvements when using in-context learning since the model is unable to learn the expected style of the long form text just from an instruction or a few examples. We sample several cases from FeTaQA as shown in Figure 5 where the ROUGE metrics assign very low scores; however, upon review, we observe that the generated answers were correct.
@@
Prompting FeTaQA
BLEU ROUGE-1 ROUGE-2 ROUGE-L
End-to-End QA 28.37 0.63 0.41 0.53
Dater (Ye et al., 2023) 29.47 0.63 0.41 0.53
@@ Appendix D Tabular Format Encoding Comparison
In alignment with prior studies Liu et al. (2023b, 2021); Jiang et al. (2022) and the baseline methods Cheng et al. (2022); Ye et al. (2023), we adopt PIPE encoding in Chain-of-Table (as shown in Appendix E). This decouples the performance gains of the proposed tabular CoT with atomic operations from the influence of various table formatting choices.
To further understand the impact of different encoding methods on table understanding performance, we conduct additional experiments using 3 additional table representations: HTML, TSV, and Markdown. For these experiments, we use End-to-End QA on WikiTQ with PaLM 2 as a running example. The results are shown in Table 10.
@@
These findings show that different tabular format encoding methods lead to different outcomes. Notably, the PIPE format adopted in our study yields the highest performance among the four encoding methods tested.

### 19 Jan 2024 v2

@@ 1 Introduction
On the other hand, inference on tables typically involves a series of intermediate reasoning steps and each of them aligns with specific tabular operations. We propose Chain-of-Table, where we conduct step-by-step reasoning as step-by-step tabular operations to form a chain of tables. The tables in the chain are the transformed tables by the tabular operations, representing the intermediate reasoning results. This procedure resembles the thought of reasoning in Chain-of-Thought (Wei et al., 2022). Specifically, we define a set of table operations, such as adding columns, selecting rows, grouping, and more, which are commonly-used in SQL and DataFrame development  (Pönighaus, 1995; Shi et al., 2020; Katsogiannis-Meimarakis & Koutrika, 2023). We then prompt LLMs to conduct step-by-step reasoning. In each step, the LLM dynamically generates an operation as the next step along with its required arguments, and then we execute the operation on the table programmatically. This operation can either enrich the table by adding detailed intermediate results or condense it by removing irrelevant information. Intuitively, visualizing the intermediate results is essential for reaching correct predictions. We feed the transformed table back for the next step. This iterative process continues until an ending state is achieved. We argue that the tables obtained during the reasoning steps are better structured representations of the intermediate thoughts than free-form text. Finally, the Chain-of-Table reasoning results in tables from which it is easier for LLMs to derive a final answer to the question.
We validate Chain-of-Table with three tabular benchmarks to evaluate table-based reasoning: WikiTQ (Pasupat & Liang, 2015), TabFact (Chen et al., 2019), and FeTaQA (Nan et al., 2022). We conduct our experiments using proprietary PaLM 2 (Anil et al., 2023) and GPT-3.5 (Brown et al., 2020; OpenAI, 2023), and open-sourced LLaMA 2 (Touvron et al., 2023), to demonstrate that our proposed method Chain-of-Table is able to generalize to various LLM options. We summarize our contribution as follows:
@@ 2 Related Work
LLMs can learn from a few samples as prompts through in-context learning. This strategy is widely used to give models additional instructions to better solve downstream tasks. Chain-of-Thought (CoT) (Wei et al., 2022) proposes to generate reasoning steps before answering instead of directly generating an end-to-end answer. Following CoT, Least-to-Most (Zhou et al., 2022) and DecomP (Khot et al., 2022) propose to break down the question into subproblems in the reasoning chain. During reasoning, the latter steps are aware of the previous ones. Such iterative chains with task decomposition further improve the results on complex problems by leveraging the intermediate results from solving subproblems. Jin & Lu (2023) enhances CoT through a table-filling procedure, with a primary focus on text-based tasks where the input and output are in textual format. However, the line of works following CoT is not specifically designed for tabular data. As reported in Chen (2023), large language models with these generic reasoning methods can achieve decent results, but there are still gaps between these methods and those specialized for table scenarios (Cheng et al., 2022; Ye et al., 2023). We propose Chain-of-Table to fill the gap by directly incorporating intermediate tables from tabular operations as a proxy of intermediate thoughts.
To better solve table-based tasks with LLMs, researchers go beyond general text and resort to using external tools. Chen et al. (2022); Gao et al. (2023) propose solving reasoning tasks by generating Python programs, which are then executed using the Python interpreter. This approach greatly improves the performance of arithmetic reasoning. In the scenario of table understanding, Text-to-SQL with LLMs (Rajkumar et al., 2022) is a straightforward application of this idea. To further push the limits of programs, Binder (Cheng et al., 2022) generates SQL or Python programs and extends their capabilities by calling LLMs as APIs in the programs. LEVER (Ni et al., 2023) also proposes solving the table-based tasks with programs but with the additional step of verifying the generated programs with their execution results. However, the assistant programs in these program-aided methods still fall short in solving difficult cases that involve complex tables. These limitations are primarily due to the constraints of the single-pass generation process, where the LLMs lack the capability to modify the table in response to a specific question, requiring them to perform reasoning over a static table. Our method, on the contrary, is a multi-step reasoning framework that conducts tabular reasoning step by step. It transforms the tables tailored to the given question.
@@ 1 Function Chain-of-Table ($T$, $Q$):
         $\triangleright$ the beginning tag, and $\phi$ means it requires no arguments
2       repeat
@@
               $\triangleright$ the current operation chain
@@ 4 Experiments
We evaluate the proposed Chain-of-Table on three public table understanding benchmarks: WikiTQ (Pasupat & Liang, 2015), FeTaQA (Nan et al., 2022), and TabFact (Chen et al., 2019). WikiTQ and FeTaQA are datasets focused on table-based question answering. They require complex tabular reasoning over the provided table to answer questions. WikiTQ typically requires short text span answers, whereas FeTaQA demands longer, free-form responses. TabFact, on the other hand, is a table-based binary fact verification benchmark. The task is to ascertain the truthfulness of a given statement based on the table. For WikiTQ evaluation, we use the official denotation accuracy (Pasupat & Liang, 2015), and for TabFact, we employ the binary classification accuracy. Given the nature of FeTaQA, which involves comparing predictions with longer target texts, we utilize BLEU (Papineni et al., 2002), ROUGE-1, ROUGE-2, and ROUGE-L (Lin, 2004) for assessment. In our experiments, we use PaLM 2-S [^1] https://cloud.google.com/vertex-ai/docs/generative-ai/learn/generative-ai-studio, GPT 3.5 (turbo-16k-0613) [^2] http://openai.com/api/, and LLaMA 2 (Llama-2-17B-chat) [^3] https://ai.meta.com/llama/ as the backbone LLMs. We incorporate few-shot demo samples from the training set into the prompts to perform in-context learning. Examples of these prompts can be found in Appendix E. Details regarding the LLM inference parameters and the number of demonstration samples used are provided in Appendix C.
@@ 4.1 Baselines
End-to-End QA guides the LLM to directly produce the answer when provided with a table and a question as input prompts. Few-Shot QA operates similarly, but it includes few-shot examples of (Table, Question, Answer) triplets in the prompt, as detailed in Brown et al. (2020). We select these examples from the training set, and the model also outputs the answer directly. Chain-of-Thought (Wei et al., 2022) prompts the LLM to articulate its reasoning process in text format before delivering the question. See Appendix F for the prompts of baselines.
@@ 4.2 Results
We compare Chain-of-Table with generic reasoning methods and program-aided reasoning methods on three datasets: WikiTQ, TabFact, and FeTaQA. The results on WikiTQ and TabFact are presented in Table 1. We have additional results on FeTaQA in Appendix B. We follow the previous works and report the performance using the official evaluation pipeline [^4] Dater Ye et al. (2023) with OpenAI Codex LLM achieves 65.9% and 85.6% accuracy on WikiTQ and TabFact, respectively. It also achieves 27.96 in BLEU, 0.62 in ROUGE-1, 0.40 in ROUGE-2, and 0.52 in ROUGE-L on FeTaQA. However, because Codex is no longer publicly available, we do not compare Chain-of-Table with Dater with Codex..
@@
We also observe that our proposed Chain-of-Table is effective across all backbone models experimented, while other competing methods, such as Binder, perform better on larger LLMs but its performance decreases with smaller LLaMA 2 (Llama-2-17B-chat). We attribute this decline to Binder’s single-pass generation process. While Binder does incorporate API calls within its framework, it lacks the capability to modify and observe the transformed tables. Consequently, Binder can only perform the tabular reasoning over a static table, making it challenging to solve complicated cases with smaller LLMs.
Figure 3: Performance of Chain-of-Thought, Dater, and the proposed Chain-of-Table on WikiTQ for questions that require an operation chain of varying lengths. Our proposed atomic operations allow our proposed method Chain-of-Table to dynamically transform the input table through multiple reasoning iterations. This significantly improves performance over generic and program-aided reasoning counterparts.
Dataset Length of operation chain
1 2 3 4 5
WikiTQ 95 1308 1481 1084 341
TabFact 4 547 732 517 223
4.3 Performance Analysis under Different Operation Chain Lengths
@@
4.4 Performance Analysis under Different Table Sizes
Large tables present significant challenges to LLMs since LLMs often struggle to interpret and integrate contexts in long input prompts (Liu et al., 2023a; Ye et al., 2023). To assess the performance on tables of various sizes, we categorize the input tables from WikiTQ into 3 groups based on token count: small ($<$2000 tokens), medium (2000 to 4000 tokens) and large ($>$4000 tokens). We then compare Chain-of-Table with Dater (Ye et al., 2023) and Binder (Cheng et al., 2022), the two latest and strongest baselines, as representative methods. Detailed results are presented in Table 3.
Prompting Total # of # of generated samples
generated samples in each steps
Binder (Cheng et al., 2022) 50 Generate Neural-SQL: 50
Dater (Ye et al., 2023) 100 Decompose Table: 40; Generate Cloze: 20;
Generate SQL: 20; Query: 20
Chain-of-Table (ours) $\leq$25 DynamicPlan: $\leq$5; GenerateArgs: $\leq$19;
Query: 1
4.5 Efficiency Analysis of Chain-of-Table
We analyze the efficiency of Chain-of-Table by evaluating the number of required generated samples. We compare Chain-of-Table with Binder (Cheng et al., 2022) and Dater (Ye et al., 2023), the two latest and most competitive baseline method. The analysis results on WikiTQ are presented in Table 4. Binder generates Neural-SQL queries, requiring 50 samples for self-consistent results. Dater involves multiple delicate yet fixed steps, such as decomposing the tables and generating cloze queries for the questions. In each step, Dater also employs self-consistency to improve accuracy of the LLM outputs, leading to a high number of required generated samples. For a detailed description of these frameworks, please refer to the corresponding papers, Ye et al. (2023) and Cheng et al. (2022).
@@ 5 Conclusion
6 Reproducibility Statement
We include the prompt examples of DynamicPlan($T$,$Q$,chain) in Appendix E.1, the demo examples of GenerateArgs($T$,$Q$,f) in Appendix E.2, the prompt examples of Query($T$,$Q$) in Appendix E.3. We run the generic reasoning methods (End-to-End QA, FewShot QA, Chain-of-Thought) using the prompts reported in Appendix F. We run Text-to-SQL and Binder using the official open-sourced code and prompts in https://github.com/HKUNLP/Binder. We run Dater using the official open-sourced code and prompts in https://github.com/AlibabaResearch/DAMO-ConvAI. We revise the code to use publicly available GPT 3.5, PaLM 2, and LLaMA 2 (Section 4) as the LLM backbone instead of the OpenAI Codex due to its inaccessibility.
@@ A.2 Ablation Study
To demonstrate the effectiveness of our proposed atomic operations, we perform an ablation study by creating five leave-one-out variants of our method, each of which removes one of the pre-defined operations from the pre-defined operation pool. For example, w/o f_add_column() means f_add_column() is removed from the operation pool. As a result, the LLM is only able to plan from the remaining four operations (f_select_column, f_select_row, f_group_by, and f_sort_by) to construct operation chains. We report the results of the ablation study in Table 5.
@@ Appendix B Experiments of Chain-of-Table on FeTaQA
Table 6 shows that Chain-of-Table also improves the performance of free-form question answering on FeTaQA across all metrics, whereas Dater (Ye et al., 2023) fails to improve the ROUGE scores compared with End-to-End QA. We also observe the marginal improvement of Chain-of-Table compared with the baseline methods. We attribute this to the nature of the n-gram text similarity metrics of ROUGE-1/2/L (Lin, 2004). As discussed in Maynez et al. (2023); Dhingra et al. (2019), these metrics are known to be insensitive to capturing improvements when using in-context learning since the model is unable to learn the expected style of the long form text just from an instruction or a few examples. We sample several cases from FeTaQA as shown in Figure 5 where the ROUGE metrics assign low scores; however, upon review, we observe that the generated answers were correct.
Prompting FeTaQA
BLEU ROUGE-1 ROUGE-2 ROUGE-L
End-to-End QA 28.37 0.63 0.41 0.53
Dater (Ye et al., 2023) 29.47 0.63 0.41 0.53
@@ Appendix D Tabular Format Encoding Comparison
In alignment with prior studies Liu et al. (2023b; 2021); Jiang et al. (2022) and the baseline methods Cheng et al. (2022); Ye et al. (2023), we adopt PIPE encoding in Chain-of-Table (as shown in Appendix E). This decouples the performance gains of the proposed tabular CoT with atomic operations from the influence of various table formatting choices.
To further understand the impact of different encoding methods on table understanding performance, we conduct additional experiments using 3 additional table representations: HTML, TSV, and Markdown. For these experiments, we use End-to-End QA on WikiTQ with PaLM 2 as a running example. The results are shown in Table 10. These findings show that different tabular format encoding methods lead to different outcomes. Notably, the PIPE format adopted in our study yields the highest performance among the four encoding methods tested.